# VERSA: A Versatile Evaluation Toolkit for Speech, Audio, and Music

### v1

@@ Abstract
In this work, we introduce VERSA, a unified and standardized evaluation toolkit designed for various speech, audio, and music signals. The toolkit features a Pythonic interface with flexible configuration and dependency control, making it user-friendly and efficient. With full installation, VERSA offers 63 metrics with 711 metric variations based on different configurations. These metrics encompass evaluations utilizing diverse external resources, including matching and non-matching reference audio, text transcriptions, and text captions. As a lightweight yet comprehensive toolkit, VERSA is versatile to support the evaluation of a wide range of downstream scenarios. To demonstrate its capabilities, this work highlights example use cases for VERSA, including audio coding, speech synthesis, speech enhancement, singing synthesis, and music generation. The toolkit is available at https://github.com/shinjiwlab/versa.
@@ 1 Introduction
Extended from its prior version in Shi et al. (2024), this work introduces a versatile evaluation toolkit for speech, audio, and music: VERSA. Built on a Pythonic interface, VERSA integrates 63 metrics and more than 711 variants, offering a wide array of automatic evaluation tools tailored for speech, audio, and music. By providing diverse metrics, VERSA aims to serve as a one-stop solution for multi-domain, multi-level, and multi-focus sound evaluation across various downstream applications. With examples showcasing the use of VERSA in ESPnet-Codec and ESPnet-SpeechLM, we anticipate that the VERSA toolkit will become a key component in advancing sound generation benchmarks, addressing challenges in speech generation, and supporting multi-modal generative frameworks. The codebase is publicly available at https://github.com/shinjiwlab/versa. [^2] A video demonstration is available at https://youtu.be/t7UP1uFvaCM.
@@ 2.1 VERSA Framework
Table 1: List of supported metrics in VERSA. The “Base" column suggests whether the metrics are included in the minimum installation. The “Model Based” column represents metrics that need pre-trained models.
No. Type Domain Name Base Variants Range Model Based Reference
@@
1 Independent Deep Noise Suppression MOS Score of P.835 (DNSMOS P.835) 1 [1, 5] Reddy et al. (2022)
2 Deep Noise Suppression MOS Score of P.808 (DNSMOS P.808) 1 [1, 5] Reddy et al. (2021)
3 Speech Quality and Naturalness Assessment (NISQA) 3 [1, 5] Mittag et al. (2021)
4 UTokyo-SaruLab System for VoiceMOS 2022 (UTMOS) 1 [1, 5] Saeki et al. (2022)
5 Packet Loss Concealment-focus MOS (PLCMOS) 1 [1, 5] Diener et al. (2023)
6 Torch-Squim PESQ (TS-PESQ) 1 [1, 5] Kumar et al. (2023)
7 Torch-Squim STOI (TS-STOI) 1 [0, 1] Kumar et al. (2023)
8 Torch-Squim SI-SNR (TS-SI-SNR) 1 (-inf, inf) Kumar et al. (2023)
9 Singing voice MOS (SingMOS) 1 [1, 5] Tang et al. (2024)
@@
10 Subjective Speech Quality Assessment (SSQA) in SHEET Toolkit 1 [1, 5] Huang et al. (2024a)
11 UTokyo-SaruLab System for VoiceMOS 2024 (UTMOSv2) 1 [1, 5] Baba et al. (2024)
@@
12 Speech Quality with Contrastive Regression (SCOREQ) wo. Ref. 2 [1, 5] Ragano et al. (2024b)
13 Speech Enhancement-based SI-SNR (SE-SI-SNR) 69 (-inf, inf) Zhang et al. (2024a)
14 Speech Enhancement-based CI-SDR (SE-CI-SDR) 69 (-inf, inf) Zhang et al. (2024a)
15 Speech Enhancement-based SAR (SE-SAR) 69 (-inf, inf) Zhang et al. (2024a)
16 Speech Enhancement-based SDR (SE-SDR) 69 (-inf, inf) Zhang et al. (2024a)
17 Prompting Audio-Language Models (PAM) metric 1 [0, 1] Deshmukh et al. (2024)
@@
18 Speech-to-reverberation Modulation Energy Ratio (SRMR) 1 [0, inf) Falk et al. (2010)
19 Voice Activity Detection (VAD) 1 - SileroTeam (2024)
20 Speaker Turn Taking (SPK-TT) 1 [0, inf) Goodwin and Heritage (1990)
21 Speaking Word Rate (SWR) 7 [0, inf) Radford et al. (2023)
22 Anti-spoofing Score (SpoofS) 2 [0, 1] Jung et al. (2021)
23 Dependent Mel Cepstral Distortion (MCD) 1 [0, inf) Kubichek (1993)
24 F0 Correlation (F0-CORR) 1 [-1, 1] Hayashi et al. (2021)
25 F0 Root Mean Square Error (F0-RMSE) 1 [0, inf) Hayashi et al. (2021)
26 Signal-to-interference Ratio (SIR) 1 (-inf, inf) Févotte et al. (2005)
27 Signal-to-artifact Ratio (SAR) 1 (-inf, inf) Févotte et al. (2005)
28 Signal-to-distortion Ratio (SDR) 1 (-inf, inf) Févotte et al. (2005)
@@ 23 Dependent Mel Cepstral Distortion (MCD) 1 [0, inf) Kubichek (1993)
29 Convolutional scale-invariant signal-to-distortion ratio (CI-SDR) 1 (-inf, inf) Boeddeker et al. (2021)
30 Scale-invariant signal-to-noise ratio (SI-SNR) 1 (-inf, inf) Luo and Mesgarani (2018)
31 Perceptual Evaluation of Speech Quality (PESQ) 1 [1, 5] Rix et al. (2001)
32 Short-Time Objective Intelligibility (STOI) 1 [0, 1] Taal et al. (2011)
33 Speech BERT Score (D-BERT) 1 [-1, 1] Saeki et al. (2024)
34 Discrete Speech BLEU Score (D-BLEU) 1 [0, 1] Saeki et al. (2024)
35 Discrete Speech Token Edit Distance (D-Distance) 1 [0, 1] Saeki et al. (2024)
36 Dynamic Time Warping Cost (WARP-Q) 1 [0, inf) Jassim et al. (2022)
@@ 30 Scale-invariant signal-to-noise ratio (SI-SNR) 1 (-inf, inf) Luo and Mesgarani (2018)
37 Speech Quality with Contrastive Regression (SCOREQ) w. Ref. 2 [1, 5] Ragano et al. (2024b)
38 2f-Model 1 [0, 100] Kastner and Herre (2019)
39 Log-weighted Mean Square Error (Log-WMSE) 1 (-inf, inf) Jordal (2023)
40 ASR-oriented Mismatch Error Rate (ASR-Mismatch) 7 [0, inf] Radford et al. (2023)
@@ 39 Log-weighted Mean Square Error (Log-WMSE) 1 (-inf, inf) Jordal (2023)
41 Virtual Speech Quality Objective Listener (VISQOL) 4 [1,5] Chinen et al. (2020)
42 Frequency-Weighted SEGmental SNR (FWSEGSNR) 1 (-inf, inf) Tribolet et al. (1978)
43 Weighted Spectral Slope (WSS) 1 [0, inf) Klatt (1982)
44 Cepstrum Distance (CD) 1 [0, inf) Barnwell III et al. (1988)
@@ 43 Weighted Spectral Slope (WSS) 1 [0, inf) Klatt (1982)
45 Composite Objective Speech Quality (Csig, Cbak, Covl) 1 [1, 5] Hu and Loizou (2007)
@@
46 Coherence and Speech Intelligibility Index (CSII) 1 [0, 1] Kates and Arehart (2005)
47 Normalized-Covariance Measure (NCM) 1 [0, 1] Chen and Loizou (2010)
@@
48 Non-match Non-matching Reference based Speech Quality Assessment (Noresqa) 2 [1, 5] Manocha et al. (2021)
49 Torch-Squim MOS (TS-MOS) 1 [1, 5] Kumar et al. (2023)
50 ESPnet ASR Model Word Error Rate (ESPnet-WER) 270 [0, inf) Watanabe et al. (2018)
@@
51 ESPnet OWSM Model Word Error Rate (OWSM-WER) 17 [0, inf) Peng et al. (2023)
@@
52 OpenAI Whisper Model Word Error Rate (Whisper-WER) 7 [0, inf) Radford et al. (2023)
53 Emotion Similarity (EMO-SIM) 1 [-1, 1] Ma et al. (2024)
54 Speaker Similarity (SPK-SIM) 15 [-1, 1] weon Jung et al. (2024)
@@
55 Non-Matching Reference Audio Quality Assessment (NOMAD) 1 [1, 5] Ragano et al. (2024a)
@@
56 Contrastive Language-Audio Pretraining Score (CLAP Score) 1 [-1, 1] Huang et al. (2023)
57 Accompaniment Prompt Adherence (APA) 1 [-1, 1] Grachten (2024)
58 Log Likelihood Ratio (LLR) 1 [0, inf) Hu and Loizou (2007)
59 Distributional Fréchet Audio Distance Audio Distance (FAD) 11 [0, inf) Roblek et al. (2019)
60 Kullback-Leibler Divergence on Embedding Distribution (KLD) 11 [0, inf) Roblek et al. (2019)
61 Density in Embedding Space (Density-Embedding) 11 [0, inf) Naeem et al. (2020)
62 Coverage in Embedding Space (Coverage-Embedding) 11 [0, 1] Naeem et al. (2020)
63 Kernel Distance/Maximum Mean Discrepancy (KID) 11 [0, inf) Roblek et al. (2019)
- Total 54 22 22 - 39 711 - 46 -
@@ 58 Log Likelihood Ratio (LLR) 1 [0, inf) Hu and Loizou (2007)
VERSA https://github.com/shinjiwlab/versa 63
@@ 2.2 Supported Metrics
As summarized in Table 1, VERSA supports a total of 63 metrics, 39 of which are included in the minimal installation. Of these, 54 metrics are applicable to speech tasks, 22 to general audio tasks, and 22 to music tasks. Additionally, several metrics feature variations based on different pre-trained models, such as word error rate, speaker similarity, and Fréchet Audio Distance (FAD) scores. By simply modifying the configuration file, [^3] An example of a configuration change is provided in Appendix A.2. VERSA can generate up to 711 distinct metric variants, offering flexibility for a wide range of evaluation scenarios.
@@ 3 Comparison to Other Toolkits
A summary of the related toolkits is presented in Table 2. Each framework has made significant contributions to the community, many serving as foundational tools for sound generation research. Building on these previous toolkits, VERSA distinguishes itself with a general design applicable across multiple domains and its comprehensive inclusion of 63 metrics with 711 variants—capabilities that have not been achieved before.
@@ 4 Demonstration
Speech enhancement: we also demonstrate the VERSA usage in speech enhancement in three speech enhancement models. More details and corresponding are shown in Table 7, Appendix C.3.
Singing synthesis is an intersection between speech and music generation, where both speech-oriented and music-oriented evaluation metrics are needed. VERSA offers a one-stop solution to this problem based on the collection of a variety of metrics in each domain. In this work, we demonstrate the evaluation of a range of singing voice synthesis models. More details and corresponding are shown in Table 9, Appendix C.4.
Music generation: the evaluation of music generation systems has received increasing attention due to the rapid progress in model development. To consider both creativity and musical harmony, recent evaluation methods mostly utilize distribution metrics for the evaluation, exhibiting a large difference to other sound generation systems. In this demonstration, VERSA aggregates a range of music generation evaluation metrics into a single toolkit. More details and corresponding are shown in Table 11, Appendix C.5.

### v2

@@ Abstract
In this work, we introduce VERSA, a unified and standardized evaluation toolkit designed for various speech, audio, and music signals. The toolkit features a Pythonic interface with flexible configuration and dependency control, making it user-friendly and efficient. With full installation, VERSA offers 65 metrics with 729 metric variations based on different configurations. These metrics encompass evaluations utilizing diverse external resources, including matching and non-matching reference audio, text transcriptions, and text captions. As a lightweight yet comprehensive toolkit, VERSA is versatile to support the evaluation of a wide range of downstream scenarios. To demonstrate its capabilities, this work highlights example use cases for VERSA, including audio coding, speech synthesis, speech enhancement, singing synthesis, and music generation. The toolkit is available at https://github.com/wavlab-speech/versa.
@@ 1 Introduction
Extended from its prior version in Shi et al. (2024), this work introduces a versatile evaluation toolkit for speech, audio, and music: VERSA. Built on a Pythonic interface, VERSA integrates 65 metrics and more than 729 variants, offering a wide array of automatic evaluation tools tailored for speech, audio, and music. By providing diverse metrics, VERSA aims to serve as a one-stop solution for multi-domain, multi-level, and multi-focus sound evaluation across various downstream applications. With examples showcasing the use of VERSA in ESPnet-Codec and ESPnet-SpeechLM, we anticipate that the VERSA toolkit will become a key component in advancing sound generation benchmarks, addressing challenges in speech generation, and supporting multi-modal generative frameworks. The codebase is publicly available at https://github.com/wavlab-speech/versa. [^2] A video demonstration is available at https://youtu.be/t7UP1uFvaCM.
@@ 2.1 VERSA Framework
Table 1: List of supported metrics in VERSA. The “Base" column indicates whether the metrics are included in the minimum installation of VERSA. The “Model Based” column represents metrics that need pre-trained models. The “Target Direction” column indicates which direction is desirable for each metric without being overly technical.
No. Type Domain Name Base Variants Range Model Based Target Direction Reference
@@
1 Independent Deep Noise Suppression MOS Score of P.835 (DNSMOS P.835) 1 [1, 5] $\uparrow$ Reddy et al. (2022)
2 Deep Noise Suppression MOS Score of P.808 (DNSMOS P.808) 1 [1, 5] $\uparrow$ Reddy et al. (2021)
3 Speech Quality and Naturalness Assessment (NISQA) 3 [1, 5] $\uparrow$ Mittag et al. (2021)
4 UTokyo-SaruLab System for VoiceMOS 2022 (UTMOS) 1 [1, 5] $\uparrow$ Saeki et al. (2022)
5 Packet Loss Concealment-focus MOS (PLCMOS) 1 [1, 5] $\uparrow$ Diener et al. (2023)
6 Torch-Squim PESQ (TS-PESQ) 1 [1, 5] $\uparrow$ Kumar et al. (2023)
7 Torch-Squim STOI (TS-STOI) 1 [0, 1] $\uparrow$ Kumar et al. (2023)
8 Torch-Squim SI-SNR (TS-SI-SNR) 1 (-inf, inf) $\uparrow$ Kumar et al. (2023)
9 Singing voice MOS (SingMOS) 1 [1, 5] $\uparrow$ Tang et al. (2024)
@@
10 Subjective Speech Quality Assessment (SSQA) in SHEET Toolkit 1 [1, 5] $\uparrow$ Huang et al. (2024a)
11 UTokyo-SaruLab System for VoiceMOS 2024 (UTMOSv2) 1 [1, 5] $\uparrow$ Baba et al. (2024)
@@
12 Speech Quality with Contrastive Regression (SCOREQ) wo. Ref. 2 [1, 5] $\uparrow$ Ragano et al. (2024b)
13 Speech Enhancement-based SI-SNR (SE-SI-SNR) 69 (-inf, inf) $\uparrow$ Zhang et al. (2024a)
14 Speech Enhancement-based CI-SDR (SE-CI-SDR) 69 (-inf, inf) $\uparrow$ Zhang et al. (2024a)
15 Speech Enhancement-based SAR (SE-SAR) 69 (-inf, inf) $\uparrow$ Zhang et al. (2024a)
16 Speech Enhancement-based SDR (SE-SDR) 69 (-inf, inf) $\uparrow$ Zhang et al. (2024a)
17 Prompting Audio-Language Models (PAM) metric 1 [0, 1] $\uparrow$ Deshmukh et al. (2024)
@@
18 Speech-to-reverberation Modulation Energy Ratio (SRMR) 1 [0, inf) $\uparrow$ Falk et al. (2010)
19 Voice Activity Detection (VAD) 1 - - SileroTeam (2024)
20 Speaker Turn Taking (SPK-TT) 1 [0, inf) - Goodwin and Heritage (1990)
21 Speaking Word Rate (SWR) 7 [0, inf) - Radford et al. (2023)
22 Anti-spoofing Score (SpoofS) 2 [0, 1] $\uparrow$ Jung et al. (2021)
23 Language Identification (LID) 17 - - Peng et al. (2023)
24 Audiobox Aesthetics (AA) 1 [1, 10] $\uparrow$ Tjandra et al. (2025)
25 Dependent Mel Cepstral Distortion (MCD) 1 [0, inf) $\downarrow$ Kubichek (1993)
26 F0 Correlation (F0-CORR) 1 [-1, 1] $\uparrow$ Hayashi et al. (2021)
27 F0 Root Mean Square Error (F0-RMSE) 1 [0, inf) $\downarrow$ Hayashi et al. (2021)
28 Signal-to-interference Ratio (SIR) 1 (-inf, inf) $\uparrow$ Févotte et al. (2005)
29 Signal-to-artifact Ratio (SAR) 1 (-inf, inf) $\uparrow$ Févotte et al. (2005)
30 Signal-to-distortion Ratio (SDR) 1 (-inf, inf) $\uparrow$ Févotte et al. (2005)
@@ 25 Dependent Mel Cepstral Distortion (MCD) 1 [0, inf) $\downarrow$ Kubichek (1993)
31 Convolutional scale-invariant signal-to-distortion ratio (CI-SDR) 1 (-inf, inf) $\uparrow$ Boeddeker et al. (2021)
32 Scale-invariant signal-to-noise ratio (SI-SNR) 1 (-inf, inf) $\uparrow$ Luo and Mesgarani (2018)
33 Perceptual Evaluation of Speech Quality (PESQ) 1 [1, 5] $\uparrow$ Rix et al. (2001)
34 Short-Time Objective Intelligibility (STOI) 1 [0, 1] $\uparrow$ Taal et al. (2011)
35 Speech BERT Score (D-BERT) 1 [-1, 1] $\uparrow$ Saeki et al. (2024)
36 Discrete Speech BLEU Score (D-BLEU) 1 [0, 1] $\uparrow$ Saeki et al. (2024)
37 Discrete Speech Token Edit Distance (D-Distance) 1 [0, 1] $\uparrow$ Saeki et al. (2024)
38 Dynamic Time Warping Cost (WARP-Q) 1 [0, inf) $\uparrow$ Jassim et al. (2022)
@@ 32 Scale-invariant signal-to-noise ratio (SI-SNR) 1 (-inf, inf) $\uparrow$ Luo and Mesgarani (2018)
39 Speech Quality with Contrastive Regression (SCOREQ) w. Ref. 2 [1, 5] $\uparrow$ Ragano et al. (2024b)
40 2f-Model 1 [0, 100] $\uparrow$ Kastner and Herre (2019)
41 Log-weighted Mean Square Error (Log-WMSE) 1 (-inf, inf) $\uparrow$ Jordal (2023)
42 ASR-oriented Mismatch Error Rate (ASR-Mismatch) 7 [0, inf] $\downarrow$ Radford et al. (2023)
@@ 41 Log-weighted Mean Square Error (Log-WMSE) 1 (-inf, inf) $\uparrow$ Jordal (2023)
43 Virtual Speech Quality Objective Listener (VISQOL) 4 [1,5] $\uparrow$ Chinen et al. (2020)
44 Frequency-Weighted SEGmental SNR (FWSEGSNR) 1 (-inf, inf) $\uparrow$ Tribolet et al. (1978)
45 Weighted Spectral Slope (WSS) 1 [0, inf) $\downarrow$ Klatt (1982)
46 Cepstrum Distance (CD) 1 [0, inf) $\downarrow$ Barnwell III et al. (1988)
@@ 45 Weighted Spectral Slope (WSS) 1 [0, inf) $\downarrow$ Klatt (1982)
47 Composite Objective Speech Quality (Csig, Cbak, Covl) 1 [1, 5] $\uparrow$ Hu and Loizou (2007)
@@
48 Coherence and Speech Intelligibility Index (CSII) 1 [0, 1] $\uparrow$ Kates and Arehart (2005)
49 Normalized-Covariance Measure (NCM) 1 [-1, 1] $\uparrow$ Chen and Loizou (2010)
@@
50 Non-match Non-matching Reference Speech Quality Assessment (Noresqa) 2 [1, 5] $\uparrow$ Manocha et al. (2021)
51 Torch-Squim MOS (TS-MOS) 1 [1, 5] $\uparrow$ Kumar et al. (2023)
52 ESPnet ASR Model Word Error Rate (ESPnet-WER) 270 [0, inf) $\downarrow$ Watanabe et al. (2018)
@@
53 ESPnet OWSM Model Word Error Rate (OWSM-WER) 17 [0, inf) $\downarrow$ Peng et al. (2023)
@@
54 OpenAI Whisper Model Word Error Rate (Whisper-WER) 7 [0, inf) $\downarrow$ Radford et al. (2023)
55 Emotion Similarity (EMO-SIM) 1 [-1, 1] $\uparrow$ Ma et al. (2024)
56 Speaker Similarity (SPK-SIM) 15 [-1, 1] $\uparrow$ weon Jung et al. (2024)
@@
57 Non-Matching Reference Audio Quality Assessment (NOMAD) 1 [1, 5] $\uparrow$ Ragano et al. (2024a)
@@
58 Contrastive Language-Audio Pretraining Score (CLAP Score) 1 [-1, 1] $\uparrow$ Huang et al. (2023)
59 Accompaniment Prompt Adherence (APA) 1 [-1, 1] $\uparrow$ Grachten (2024)
60 Log Likelihood Ratio (LLR) 1 [0, inf) $\uparrow$ Hu and Loizou (2007)
61 Distributional Fréchet Audio Distance Audio Distance (FAD) 11 [0, inf) $\downarrow$ Roblek et al. (2019)
62 Kullback-Leibler Divergence on Embedding Distribution (KLD) 11 [0, inf) $\downarrow$ Roblek et al. (2019)
63 Density in Embedding Space (Density-Embedding) 11 [0, inf) $\uparrow$ Naeem et al. (2020)
64 Coverage in Embedding Space (Coverage-Embedding) 11 [0, 1] $\uparrow$ Naeem et al. (2020)
65 Kernel Distance/Maximum Mean Discrepancy (KID) 11 [0, inf) $\downarrow$ Roblek et al. (2019)
- Total 56 23 23 - 40 729 - 48 -
@@ 60 Log Likelihood Ratio (LLR) 1 [0, inf) $\uparrow$ Hu and Loizou (2007)
VERSA https://github.com/shinjiwlab/versa 65
@@ 2.2 Supported Metrics
As summarized in Table 1, VERSA supports a total of 65 metrics, 39 of which are included in the minimal installation. Of these, 54 metrics are applicable to speech tasks, 22 to general audio tasks, and 22 to music tasks. Additionally, several metrics feature variations based on different pre-trained models, such as word error rate, speaker similarity, and Fréchet Audio Distance (FAD) scores. By simply modifying the configuration file, [^3] An example of a configuration change is provided in Appendix A.2. VERSA can generate up to 729 distinct metric variants, offering flexibility for a wide range of evaluation scenarios.
@@ 3 Comparison to Other Toolkits
A summary of the related toolkits is presented in Table 2. Each framework has made significant contributions to the community, many serving as foundational tools for sound generation research. Building on these previous toolkits, VERSA distinguishes itself with a general design applicable across multiple domains and its comprehensive inclusion of 65 metrics with 729 variants—capabilities that have not been achieved before.
@@ 4 Demonstration
Speech enhancement targets the improvement of speech transmission in different environments. In this paper, we also demonstrate the VERSA usage in speech enhancement in three speech enhancement models. More details and corresponding information are shown in Table 7, Appendix C.3.
Singing synthesis is an intersection between speech and music generation, where both speech-oriented and music-oriented evaluation metrics are needed. VERSA offers a one-stop solution to this problem based on the collection of a variety of metrics in each domain. In this work, we demonstrate the evaluation of a range of singing voice synthesis models. Table 9, Appendix C.4 shows more details and corresponding information.
Music generation and its evaluation have received increasing attention due to the rapid progress in model development. To consider both creativity and musical harmony, recent evaluation methods mostly utilize distribution metrics for the evaluation, exhibiting a large difference to other sound generation systems. In this demonstration, VERSA aggregates a range of music generation evaluation metrics into a single toolkit. More details and corresponding information are shown in Table 11, Appendix C.5.